OpenStudy (anonymous):

if f(x) = -x^-4x+3 find f(-2)

OpenStudy (anonymous):

f(-2) = -2^(-4*-2+3) // - I am assuming f(x) = 0x^(-4x+3) f(-2) = -2^(8+3) =-2^11 =-2048

OpenStudy (anonymous):

\[f(x)=-x ^{2}-4x+3\]\[f(-2)=-(-2) ^{2}-4(-2)+3\]\[f(-2)=-(4) +8+3\] f(-2)=7

OpenStudy (anonymous):

yes that ansewr is correct